### v1

# A Blue Start: A large-scale pairwise and higher-order social network dataset

Abstract
Large-scale networks have been instrumental in shaping the way that we think about how individuals interact with one another, developing key insights in mathematical epidemiology, computational social science, and biology. However, many of the underlying social systems through which diseases spread, information disseminates, and individuals interact are inherently mediated through groups of arbitrary size, known as higher-order interactions. There is a gap between higher-order dynamics of group formation and fragmentation, contagion spread, and social influence and the data necessary to validate these higher-order mechanisms. Similarly, few datasets bridge the gap between these pairwise and higher-order network data. Because of its open API, the Bluesky social media platform provides a laboratory for observing social ties at scale. In addition to pairwise following relationships, unlike many other social networks, Bluesky features user-curated lists known as ”starter packs” as a mechanism for social network growth. We introduce ”A Blue Start”, a large-scale network dataset comprising 26.7M users and their 1.6B pairwise following relationships and 301.3K groups representing starter packs. This dataset will be an essential resource for the study of higher-order network science.
I Background & Summary
Network science has revolutionized the study of human behavior because of its ability to represent complex relationships between individuals in a population. Representing social structures as networks offers insights on the temporal nature of interactions [1], the heterogeneity of interaction patterns [2, 3, 4], preferential relationships between nodes with similar characteristics [5, 6] and highly connected nodes [7], and community structure [8]. One particular medium that spawned an entire sub-field of network science was Twitter, the social media platform. This online ”laboratory” was particularly useful because of its open automated programming interface (API); the free ”Twitter Streaming API” provided researchers with up to 1% of all tweets, while the paid ”Twitter Firehose” allowed researchers to extract 100% of all public posts [9]. As Twitter grew in size, the Firehose API was discontinued and replaced with the Decahose API which allowed researchers to download 10% of all posts in real time [10, 11]. Nevertheless, the relative availability of Twitter data, along with the platform’s popularity and use across demographic and sociographic groups [12] spurred scientific innovations across a number of disciplines, including epidemic modeling [13], political communication [14], and statistical physics [15], to name just a few. Some sub-disciplines that depend on access to ”everyday” communication, such as social movement studies, were revolutionized by access to the Twitter API, which provided data to support hundreds of papers on the structure and discursive content of social movements around the world (e.g., see Refs. [16, 17, 18]).
The change in ownership of Twitter, subsequently known as X, spawned interest in many alternative platforms such as Mastodon, Threads, and Bluesky [19]. We focus our attention of Bluesky because of its inherent open access to platform data. Bluesky originated in 2019 as a research initiative at Twitter examining the potential for open-sourcing and decentralizing social media [20]. It was incorporated as an independent company in 2021  [21]. Bluesky’s central innovation is a standard known as the Authenticated Transfer (AT) Protocol [22]. The AT Protocol emphasizes federated identities, allowing users to own and curate a single online identity that is usable on all AT Protocol platforms and services [23]. User data is stored in repositories on personal data servers (PDSs) that log user actions; these repositories are intended to be portable so that users can easily migrate between hosts [24]. The Bluesky application queries a user’s PDS to deliver content to users; however, anyone can build an application to view AT Protocol data. Applications with access to the same data can vary in their content moderation policies, post visibility, or algorithmic curation strategies, for example [25].
Although large-scale social network datasets derived from online social media platforms such as Twitter [9], Bluesky [26, 27, 28, 29], Facebook [30], and many others [31, 32, 33, 34, 35, 36, 37, 38, 39] have formed the experimental basis of many studies on human behavior, these studies almost exclusively analyze these networks from a pairwise perspective, since following relationships are inherently pairwise. In contrast, many social systems contain interactions which may not be pairwise but rather may involve an arbitrary number of individuals. Higher-order networks are composed of the collection of these interactions, known as higher-order interactions, and can more accurately represent the structure of many empirical systems [40, 41, 42]. For example, in the scientific ecosystems, grants and scientific publications often involve large teams of collaborators, and the size [43, 44, 45], interdisciplinarity [46], and composition [47] of these teams can be linked to scientific outcomes such as citations and collaboration trajectories. However, co-authorship networks have historically been represented as pairwise networks [48], where a publication with k authors is converted to a $k$-clique, making it challenging to disambiguate, for example, a paper authored by A, B, and C, or three papers authored by A and B, B ad C, and A and C. Similarly, in computational social science, higher-order networks can help model the formation and fragmentation of groups [49], face-to-face interactions [50], and network structure at different scales of interaction [40]. There are relatively few truly higher-order social network datasets, however. Of those that do exist, they can largely be categorized as email datasets [51] and proximity datasets [52, 1, 53, 54] for which maximal simultaneous cliques of pairwise proximity measurements have been converted to higher-order interactions [51]. In fact, several statistical approaches have been developed to infer higher-order interactions from pairwise networks [51, 55, 56]. However, there is a lack of both intrinsically higher-order social network datasets and datasets comprising both higher-order interactions and pairwise interactions.
This Bluesky dataset addresses both of these gaps in the corpus of higher-order network datasets. In the summer of 2024, Bluesky developers announced a new feature in the Bluesky app, known as starter packs (SPs). Starter packs are user-curated lists of accounts and feeds (which serve post content to users based on custom algorithms) with a user-supplied title and description. Users share these lists so that other users can quickly grow the list of accounts they follow based on specified interest. One has the option to either browse the list and cherry-pick accounts to follow or hit the ”follow all” button to follow all accounts in the list. Starter packs are inherently higher-order. At the time of this writing, users can create starter packs with 8 to 150 users and feeds. There is a caveat, however, which is that a similar higher-order structure on Bluesky known as ”lists”, which are not size restricted and so contain anywhere between 0 and several thousand items. With the advent of starter packs, users developed tools for converting lists into starter packs, bypassing the standard size limits. So while the list of starter packs is almost entirely higher-order, there are exceptions to this (as can be later seen in Fig. 1). One can also extract the list of accounts that each user follows, which can be represented as a directed, pairwise network. These two types of relationships bridge the gap between pairwise and higher-order network data. Furthermore, this dataset offers researchers ground to investigate social phenomena exclusive to groups phenomena and develop measures and algorithms tailored to such structures. In particular, the absence of edges in the following network when compared to the network induced by starter packs facilitates exploration of patterns that are exclusive to higher-order networks. For example, one can analyze user overlap patterns between starter packs, a structural feature inaccessible to pairwise network analysis. Higher-order network analysis enables a deeper understanding of the complex dynamics among users participating in group interactions.
Beyond applications for the study of higher-order dynamics and interactions, our data enables novel research in communication and human-computer interaction. Future work in this area can leverage our dataset to address questions of structural differences in start pack inclusion, the effect of starter pack inclusion on following relationships, and more. These questions also lend themselves to salient new possibilities in graph machine learning. In this setting, empirical data on starter packs becomes crucial for training and testing models. More generally, the broad applicability of online social network data across scientific fields is well documented in the extensive literature that has emerged over the past decade, leveraging existing datasets from other platforms. Now, with up-to-date data becoming increasingly difficult to access, this dataset provides a unique opportunity to gain early insights into a key, distinguishing feature of a burgeoning new online social media platform.
II Methods
II.1 Obtaining the list of users
The complete list of users and all operations performed on a user’s account, such account creation, modification, and deletion, are stored in the ”DID PLC Directory” at https://plc.directory. User accounts are assigned a ”decentralized identifier” (DID) which provides a persistent global reference for all account information and all queries to the Bluesky API to collect user information such as the accounts followed or the list of starter packs require a DID as a query parameter. Although in principle one could start from a given user and their associated DID and perform a breadth-first search through the following relationships to extract the list of users, the network is disconnected and, as seen in Fig. 5(b), some of the strongly-connected components are non-negligible in size. The export endpoint enables efficient downloading of the contents of the PLC Directory and we extracted the list of unique DIDs from the exported data in batches of 1,000. Although the PLC Directory also tracks account deactivation or deletion, we compiled the list of DIDs without considering this information. We collected the list of DIDs on January 22nd, 2025.
II.2 Collecting starter packs
We used the public app.bsky.graph.getActorStarterPacks public API endpoint to obtain the list of all starter packs associated with a user account. For each unique account DID extracted from the directory mentioned above, we queried this endpoint and extracted the Uniform Resource Identifier (URI) corresponding to the list representation of the starter pack for each starter pack the user created. Next, we queried the app.bsky.graph.getList public API endpoint with each URI associated with the user account and obtained the following information on each starter pack: (1) the name and description of the starter pack; (2) the creation date; (3) the starter pack creator’s DID, handle, and account creation date; and (4) the list of user accounts and feeds comprising the starter packs. Querying the app.bsky.graph.getList endpoint returns a cursor which allowed us to page through in batches of 100 starter pack items and collect the entire list of entities in the starter pack. The name and description of a starter pack are optional; if a user does not specify a name, it defaults to ”{user display name}’s Starter Pack” and if a user does not write a description, it is left empty. Note that entities in a starter pack are not necessarily users, they can be any entity that can be ”followed” on Bluesky (e.g. a list or a feed). Lists can only be followed, however, by circumventing the standard user interface through the API. We used the Python asyncio and aiohttp packages to efficiently download data from the API while respecting the rate limits [^1] https://docs.bsky.app/docs/advanced-guides/rate-limits imposed by the Bluesky API. We downloaded data on each starter pack as a Python dictionary and appended each pack to a list of all starter packs. We then saved this list to a JSON blob, described more in Section III.1, and used the XGI Python package [58] to convert the starter pack network to the HIF standard [^2] https://github.com/pszufe/HIF-standard and a hyperedge list. We collected the list of starter packs over several days ending on February 3rd, 2025.
II.3 Collecting network data
For each account DID collected from querying the DID PLC directory, we obtained the list of users followed by that account by querying the app.bsky.graph.getFollows public API endpoint. Querying this endpoint returns a cursor which allowed us to page through batches of 100 following relationships and collate the list of all accounts followed. This process yielded a snapshot of all following relationships for all accounts corresponding to the DID PLC directory. We used asynchronous programming, as explained in Section II.2, to efficiently download the data. We collected the user following relationships over several days ending on February 3rd, 2025.
II.4 Anonymizing user data
Although all user data is available via the public Bluesky API and users can require authentication to access their public user data from the API, we nonetheless take reasonable steps to prevent re-identification of users. We recognize, however, that with enough effort, that it is possible to re-identify all users. However, we assert that this effort will be comparable to querying the API from scratch. We anonymize our dataset in three ways: (1) by converting the DIDs of all users and feeds to unique integer identifiers, (2) by removing the name and description of all starter packs, and (3) by coarsening the precision of all timestamps to one day. We anonymize the list of starter pack members and user accounts by first taking the union of the list of entities in the higher-order network of starter packs and the list of all user accounts, then assigning integers to every listed entity. We do the same for each starter pack URI, converting it to an integer. Then we iterate through every starter pack and every following relationship using this mapping from DIDs and URIs to unique integers. This process ensures that while users cannot be identified, node IDs are consistent across both the following network and the starter pack network.
II.5 Limitations
The approach we utilized in collecting the starter pack and following data comes with limitations, however. The number of active accounts, starter packs, and following relationships on the Bluesky platform is rapidly changing and any collection method not taking this into account will not perfectly capture the network structure. Using the PLC SIS Directory as a static directory does not consider accounts created after this list was collected. This leads to collected DIDs in the starter pack and following data that are not present in the ”global” list of DIDs; existing users can both follow and add to a starter pack users that have been created since the list of DIDs was collected. Secondly, time stamps and user data can be arbitrarily set by users if they interact with Bluesky directly through the AT Protocol and not through the Bluesky application. Lastly, follows (and starter pack members) can be added and removed while waiting to process a particular DID and similarly, after that DID has been processed, following relationships and curated starter packs can change as well. One way around this is by downloading streaming data from the Bluesky Firehose API while backfilling all events for all users, as is done in Ref. [29]. The downside to this approach is that it is an immense amount of data; roughly 27GB of JSON data per day at the time of writing. We believe that our approach strikes the balance of being computationally manageable, while surveying close to the entire network.
II.6 Ethical Considerations
This data collection was ruled an exempt protocol by the Institutional Review Board (IRB) at Northeastern University (No. 25-01-22) and by the Institutional Review Board for the Social and Behavioral Sciences (IRB-SBS) at the University of Virginia (No. 7332). All of the data obtained is publicly accessible (without authentication or a Bluesky account) via the Bluesky API. Additionally, this study was granted a waiver of informed consent from both the IRB at Northeastern University and the IRB-SBS at the University of Virginia; arguably, obtaining informed consent for all users’ inclusion in this dataset would constitute a larger risk to users’ privacy than the collection of publicly available data. Nonetheless, we recognize that this dataset could enable large-scale identification of accounts’ following relationships. We therefore redact all user DIDs (decentralized identifiers, used for long-term, persistent identification of entities in the AT Protocol that Bluesky uses) and truncate the timestamps on all starter pack and account creation events to include only the date; i.e., the level of precision at which timestamps are reported is one day. The details of this process are described in Section II.4.
III Data Records
This study releases two data records: (1) a list of starter packs on Bluesky and (2) a list of user following relationships.
III.1 Bluesky starter packs
We release the starter pack dataset in three different formats to ensure maximum cross-compatibility.
JSON Blob (deidentified_starterpack_blob.json, 132.7 MB). We provide a JSON file of all scraped starter packs’ contents and metadata, intended for computational analysis. This file contains a list of starter packs; each starter pack contains the fields described in Table 1.
Field Type Description
pack-id integer Anonymous integer identifier of the starter pack list
creator-id integer Anonymous integer identifier of the starter pack’s creator
creator-date-created date The date on which the SP creator’s account was created
date-created date The date on which the starter pack was created
members list of integers Each entry is a member user’s anonymous integer identifier
Table 1: Fields for the JSON Blob and the HIF-compliant files of the starter pack network.
Hyperedge list (deidentified_starterpack_edgelist.csv, 85.1 MB). We also provide a hyperedge list stored as a CSV file. Each line in this file corresponds to a starter pack, and each entry in the line is a member of the starter pack. Each line contains entries separated by commas; a single line corresponds to a single starter pack. For example, the line ”1,2,3,4” indicates that users with anonymous integer identifiers 1, 2, 3, and 4 belong to the same starter pack.
Hypergraph Interchange Format (deidentified_starterpack_hif.json, 619.1 MB). This file is a HIF-compliant JSON file. The HIF standard provides a cross-platform standard for sharing attributed higher-order network data. The edge attributes (”pack-id”, ”creator-id”, ”creator-date-created”, and ”date-created”) are labeled as described in Table 1.
III.2 Bluesky following network
The following network dataset is stored in the file deidentified_follows_edgelist.zip (8.5 GB) as a compressed CSV file. Each line $(A,B)$ in the file indicates that user $A$ follows user $B$. Here, $A$ and $B$ are anonymous integer identifiers that map to user DIDs. As described in Section II.4, these anonymous integer IDs can be linked to creators and members of starter packs.
IV Technical Validation
Here, we describe the statistical properties of both the starter pack network and the following network, demonstrating their validity and utility for researchers analyzing social network data.
IV.1 Starter pack data
The starter pack data comprises 1,671,681 unique user and feed IDs and 301,322 starter packs. These statistics corroborate the findings of Ref. [29] which assembled a dataset of 335,416 starter packs. The discrepancy in sizes between these two datasets are attributable to the time of collection, the method of collecting the dataset, and the deletion of existing starter packs. Ref. [29] collected all of the starter packs from their introduction up to, and not including, January 1st, 2025, whereas with our method, we collected the list of DIDs on January 22nd, 2025 and queried the list of starter packs over several days ending on February 3rd, 2025. Out method relied on querying the static Bluesky API, whereas Ref. [29] used the Bluesky Firehose API and the Bluesky sync API (For more details, see Sec. II.5) to collect all starter packs. The advantage of the approach in Ref. [29] is that it can collect starter packs even if they are later deleted; it is estimated that 20% of all starter packs were deleted by January 1st, 2025 [29].
Our analysis of the starter pack data is from a higher-order network perspective; particularly, we model starter packs as higher-order interactions, or hyperedges, and the collection of starter packs as a higher-order network, or hypergraph. Higher-order network analysis can uncover network structure across different scales by modeling hyperedge inclusion [60, 41], overlap [61], and interaction size [40]. Because starter packs are not dyadic, higher-order network analysis represents a natural framework to study this system. We expect this analysis to yield important insights into the group overlap structure and, more importantly, how groups and and pairwise ties can co-evolve.
Figure 1: Basic higher-order starter pack statistics. Panel (a) plots the probability distributions of both the number of starter packs to which a user account belongs and the number of starter packs that each user account has created. Panel (b) plots the starter pack size distribution, with the minimum and maximum starter pack sizes illustrated. Panel (c) plots the sizes of the connected components with the minimum and maximum starter pack sizes illustrated.
We begin our analysis by examining the distribution of the number of starter packs that users belong to and create, as seen in Fig. 1(a). The starter pack sizes range from a minimum size of 0 to a maximum size of 4,070, with a mean of 32.86, a median of 24.0, and a mode of 8, while the number of starter packs to which a user account or feed belongs varies from 1 to 145,343 (with the highest count belonging to @bsky.app), averaging 5.92 memberships with a median of 1.0 and a mode of 1. Although the number of starter packs that users create varies widely, it is also quite homogeneous; while the majority (84.2%) of accounts in the network did not create a starter pack, those who did created 1.14 starter packs on average. Fig. 1 shows the hyperedge (starter pack) size distribution in Fig. 1(b) which, interestingly, is trimodal, with peaks at 8, 50, and 150. The peaks at 8 and 150 correspond to the minimum and maximum allowed starterpack sizes, and the peak at 50 is due to the Bluesky starter pack feature called ”Make one for me,” which pre-populates a starter pack with up to 50 accounts and feeds that a user follows. The distribution of connected component sizes (Figure 1, right) shows a very large giant connected component of 1.67M users accompanied by 382 components all smaller than the starter pack size limit of 150, the vast majority (361) of which are isolated starter packs.
Figure 2: Temporal starter pack statistics. Panel (a) plots the number of starter packs created each day and panel (b) plots the distribution of time elapsed between when a Bluesky account was created and when that account created a starter pack. The dashed lines in panel (a) illustrate the following notable events: (1) Bluesky announces the release of starter packs as a feature on June 26th, 2024 [62], (2) the Brazilian government bans the X platform on August 30th, 2024, (3) changes to the blocking feature on X are announced on October 16th, 2024 allowing blocked accounts to view posts of the blocking account, and (4) X announces new terms of service on November 15th, 2024 requiring users to consent to their posts being used to train X’s artificial intelligence models.
Analyzing the times at which starter packs were created, we see four sharp peaks characterized by discontinuous jumps in the number of starter packs created and exponential decay after peak activity as illustrated in Fig. 2(a). These four peaks correspond to notable events in the history of both Bluesky and X. Much of Bluesky’s growth was spurred by user migration from X in response to platform and access changes. For example, in Fig. 2(a), we see (1) when starter packs were introduced as a feature [62], (2) when X was banned in Brazil, spurring an influx of Brazilian users, (3) the change to the blocking feature on X allowing blocked users to view the content of the account that blocked them, and (4) the change to the X terms of service, requiring users to consent to their posts being used to train X’s artificial intelligence models. Fig. 2(b) plots the distribution of time between when a user makes a Bluesky account and when they create a starter pack. This plot exhibits two interesting features: first, several starter packs are created by accounts with negative ages and second, the distribution of account ages is bimodal. Negative account ages are physically impossible; starter packs cannot be created prior to the creation of the associated user account. However, users have the ability to change their user data arbitrarily through the Bluesky API and this can cause data artifacts such as these. The distribution of account ages has a strong peak at zero and is bimodal, indicating that most users create a starter pack on the same day that they create an account and that early adopters of Bluesky are also active in creating starter packs.
Figure 3: $s$-line graph statistics of the starter pack network. The number of nodes and edges in the $s$-line graph of the starter pack network for all non-trivial values of $s$. In panels (a) and (b), the dashed line denotes the maximum starter pack size of 150.
Line graph statistics are a powerful tool for analyzing higher-order network analysis and we use them to describe the edge overlap patterns in the starter pack network. A line graph of a higher-order network is a network with nodes corresponding to hyperedges and edges connecting two nodes if the two hyperedges that they represent share at least one node in common. Ref. [61] generalized this idea by introducing $s$-line graphs which only connect two nodes with an edge if the hyperedges they represent share at least $s$ nodes in common. We vary $s$, the edge overlap size, to get an idea of the strength of overlap across different regions of the network. This can be quite computationally intensive and it is often necessary to exploit algorithms such as those described in Ref. [63]. Because our analysis was confined to measuring the number of nodes and edges in the $s$-line graph, we were able to utilize a much more memory-efficient algorithm which focused on first finding cliques in the line graph and then looking for connections between edges in different cliques. This approach allowed us to avoid storing all the edges in a clique and simply store the node labels of the clique, which reduced the memory requirements considerably. Fig. 3 plots the number of nodes and edges in each $s$-line graph. Note that these graphs are extremely large and dense; for $s=1$, 29% of all possible edges are present in the line graph. Unsurprisingly, the maximum starter pack size is the most determinant of the connectedness of the line graph, followed by the size of the ”make one for me” starter pack feature corresponding to $s=50$. This higher-order network is far more dense and overlapping than many other existing higher-order datasets, providing an exciting opportunity for studies on higher-order network formation and higher-order algorithm development.
Figure 4: Mesoscale starter pack statistics. Panel (a) plots the distribution of $k$-coreness, Panel (b) plots the probability of two user accounts appearing in $m$ different starter packs together, and Panel (c) plots the distribution of normalized entropy calculated for all starter packs. The illustrations of starter packs indicate at $s=0$, a starter pack completely contained in an inferred community and at $s=1$, a starter pack where every member belongs to a different community.
We perform a hypergraph $k$-core decomposition [64] and, as seen in Fig. 4(a), the distribution has a very heavy tail. There are over 500 users with $k$-cores of over 1000 and this is indicative of many nodes deep in the core of the hypergraph, which implies the existence of large and intricate intersections between starter packs. Fig. 4(b) plots the co-occurrence distribution of user pairs, i.e., the probability of two user accounts appearing in $m$ different starter packs together. This can also be thought of as examining the distribution of weights in a weighted clique expansion of the hypergraph. From ocular regression, the distribution appears piecewise linear, suggesting three distinct regimes. There is a substantially high chance of having very large co-occurrence numbers. These insights suggest complex co-occurrence patterns, with a plethora of users participating in a multitude of starter packs simultaneously. In future studies, it will be fruitful to examine the co-occurrence of groups of $k$ nodes across different starter packs along the lines of techniques presented in [65]. To measure the community structure of the starter pack network, we run the Leiden [66] algorithm for two iterations on the unweighted pairwise projection of the largest connected component of the hypergraph to infer the community labels of all the nodes in the starter pack network, resulting in 654 clusters. The largest connected component was used so that isolated edges and small components would not inflate the number of clusters calculated. We use the inferred community labels, $\bm{\ell}_{e}=[\ell_{1},\dots,\ell_{k}]^{T}$, for each starter pack $e$ of size $k=|e|$, to specify an associated probability mass function $p_{e}(\lambda)=\sum_{1}^{k}\bm{1}_{\ell_{k}=\lambda}/k$. Then the normalized Shannon entropy for each each starter pack is defined as $s_{e}=-\sum_{\lambda}p_{e}(\lambda)\log_{2}[p_{e}(\lambda)]/\log_{2}(k)$, where $\lambda$. By this measure, the lowest entropy configuration ($s=0$) is a starter pack composed of nodes all belonging to the same community and the highest entropy configuration ($s=1$) is a starter pack composed of nodes all belonging to different communities. As seen in Fig. 4(c), there is a small spike at $s=0$ which can be traced to starter packs with 0, 1, or 2 members. Similarly, only starter packs with thirteen or fewer members achieved the maximum entropy configuration of $s=1$. This can be attributed to the constraints induced by a finite number of community labels; as the starter pack size $k$ increases, it becomes increasingly unlikely to find an edge composed of nodes, each with a different community label. However, Fig. 4(c) indicates that the vast majority of edges are closer to the maximum entropy configuration than to the minimum entropy configuration. This seems to indicate that the starter pack hypergraph exhibits relatively little community structure.
IV.2 Following data
The following network consists of 26,692,837 distinct users and 1,654,721,704 directed following relationships linking a follower and a followed account. The minimum out-degree, or following count, in this network is 1 (a user who is not following anyone would not be included in the network), the maximum out-degree is 449,058, and the average out-degree is 64.75. The minimum in-degree, or follower count, in this network is 1, the maximum in-degree is 19,946,766, and the average in-degree is 92.47. The discrepancies in the mean degrees can be attributed to the fact that while 8,798,605 accounts have no followers, only 1,138,153 accounts do not follow any accounts. When including degree-zero nodes, the mean in- and out-degrees are both 61.99.
Figure 5: Basic following network statistics. Panel (a) plots the in-degree and out-degree distributions and panel (b) plots the frequency of strongly and weakly connected component sizes.
We begin our analysis by looking at the in- and out- degree distributions; both distributions are highly heterogeneous and heavy-tailed as seen in Fig. 5(a). In addition, we see that the in-degree distribution has a heavier tail than the out-degree distribution; not only are there more accounts following a large number of accounts, but the maximum number of accounts followed is larger than the number of followers of the most popular account. Fig. 5(b) plots the distributions of both the strongly and weakly connected components in the following network. The distribution of strongly connected component sizes contains one giant component of 16,594,500 nodes, along with several strongly connected components of 100 nodes or fewer. Looking at the distribution of weakly connected component sizes, we see an increase in the size of the giant component and a decrease in the number of small connected components, indicating that while the vast majority of the following graphs is strongly reachable, there are still ”in-components” and ”out-components” up to several hundred nodes in size. These components may or may not be reachable via snowball sampling [67] depending on the starting node of choice. A small number of components are truly disconnected from all other components via follower/following relationships.
IV.3 Comparing the starter pack and following network data
The ”A Blue Start” dataset provides a bridge between pairwise and higher-order network representations. In Fig. 6, we compare the rankings induced by degree centrality in the starter pack hypergraph and the follows network. In the starter pack hypergraph, the degree centrality $d_{s}$ of a node (user) is the number of hyperedges (starter packs) of which the user is a member, while in the following network, the degree centrality $d_{f}$ of a node (user) is the sum of its following and followed by counts. Kendall’s $\tau_{\beta}$ rank correlation coefficient between two rankings ranges from -1 (total disagreement) to 1 (total agreement) and systematically handles ties in ranking. We start by computing $d_{s}$ and $d_{f}$ and only analyzing nodes appearing in both networks. We 1) rank the top $k$ nodes according to $d_{s}$ and compute the correlation between the ranks of these same nodes according to $d_{f}$, presenting the results as $d_{s}-d_{f}$; and 2) rank the top $k$ nodes according to $d_{f}$ and compute the correlation between the ranks of these same nodes according to $d_{s}$ and present the results as $d_{f}-d_{s}$. The fact that the maximal value of 1 is achieved at $k=1$ is not surprising, as the @bsky.app account is most central in both networks. However, the rankings quickly diverge, with a weaker correlation of approximately 0.5 for $k=10$ ($d_{s}-d_{f}$); this association increases to around 0.7 at $k=100$ and decreases with increasing $k$. Overall, relatively low correlation among the importance assigned to nodes in the higher-order (starter pack) and dyadic (following network) data demonstrates that the two contain complementary information about user importance, and that importance is conditioned on the context in which user interactions are measured.
Figure 6: Comparing pairwise and higher-order network structure. Kendall’s $\tau_{\beta}$ rank correlation coefficient between degree centrality in the starter pack data ($d_{s}$) and the following data ($d_{f}$) as a function of number of top-ranked nodes compared.
V Usage Notes
Processing the following network can be computationally intensive; The CSV file contains 1.3B lines of following relationships. Simply loading this dataset using the pandas Python package [68, 69] requires roughly 25 GB of memory. The size limitation of a pandas dataframe tends to be the amount of available memory and we encourage users of this dataset to verify that they have enough physical resources to process the following network data. To parallelize the processing of these data, we recommend loading the CSV with dask [70]. We opted to convert the dataset into a nested Python dictionary when computing the component sizes. The degree distributions were computed using a groupby operation in dask, and the component sizes were computed using the py-tarjan Python package.
The starter pack data is a bit more manageable, requiring an order of magnitude less memory. However, as mentioned in Section IV.1, the line graph computed from the starter pack data is extremely dense; 29% of all possible pairs of starter packs share at least one member in common. When working with the edge connection data, one can dramatically decrease the number of shared starter pack members by removing the highest degree accounts. For example, the @bsky.app account is involved in 48% of all starter packs, and the top-5 most included accounts are all in greater than 10% of starter packs. In addition, one can decrease the number of edges by filtering out edges by their size, similar to the approach suggested in Ref. [40].
VI Data availability
All data is available on the Social Media Archive (SOMAR), which is a project by the Inter-university Consortium for Political and Social Research. The datasets described in this paper are available in the data collection titled ”A Blue Start: A large-scale pairwise and higher-order social network dataset” [71]. The dataset is available under SOMAR’s Controlled Download access method: researchers will be required to fill out a brief application detailing their research topic and agreeing to the ICPSR’s Terms of Use. This step is required in order to make sure the data are only used for research purposes; researchers must agree that they will not attempt to re-identify anonymized individuals.
The code used to analyze the starter pack and following networks is available on GitHub and at Ref. [72].
Author Contributions
A.H.S.: project conception, data acquisition, data processing, data analysis, writing. I.A.: data analysis, writing. S.K.: writing. B.F.W.: project conception, writing. N.W.L.: project conception, data acquisition, data processing, data analysis, writing.
Acknowledgements.
N.W.L. acknowledges support from the University of Virginia Prominence-to-Preeminence (P2PE) STEM Targeted Initiatives Fund, SIF176A Contagion Science. A.H.S. acknowledges support from the National Science Foundation Graduate Research Fellowship Program under grant no. 1938052. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. Pacific Northwest National Laboratory is operated by Battelle Memorial Institute under Contract DE-ACO6-76RL01830. PNNL Information Release Number: PNNL-SA-211224.